The top part of Ramakrishnan, Inc.’s 2021 and 2020 balance sheets is listed below (in millions of dollars). 2021 2020 2021 2020 Current assets: Current liabilities: Cash and marketable securities $ 52 $ 43 Accrued wages and taxes $ 41 $ 40 Accounts receivable 152 137 Accounts payable 96 85 Inventory 215 196 Notes payable 67 59 Total $ 419 $ 376 Total $ 204 $ 184 Calculate Ramakrishnan, Inc.’s current ratio for 2021 and 2020

Calculate Ramakrishnan, Inc.’s quick ratio for 2021 and 2020.

Calculate Ramakrishnan, Inc.’s cash ratio for 2021 and 2020.
Answer #1

2021

Current ratio = current assets/current liablities
Current ratio = 419/204
Current ratio = 2.05
Quick or acid test ratio = (current assets-inventory)/current liablities
Quick or acid test ratio = (419-215)/204
Quick or acid test ratio = 1
Cash ratio = cash & cash eqv./current liablities
Cash ratio = 52/204
Cash ratio = 0.25

2020

Current ratio = current assets/current liablities
Current ratio = 376/184
Current ratio = 2.04
Quick or acid test ratio = (current assets-inventory)/current liablities
Quick or acid test ratio = (376-196)/184
Quick or acid test ratio = 0.98
Cash ratio = cash & cash eqv./current liablities
Cash ratio = 43/196
Cash ratio = 0.22